OpenStudy (riley):

Find the derivative of the function. s(x) = (e^4x − 1)/(x^3 − 1)

OpenStudy (anonymous):

Product Rule: s '(x) = (e^4x-1)(3x^2) + (4e^4x)(x^3 - 1)

OpenStudy (anonymous):

\[\frac{3 x^2-e^4 \left(2 x^3+1\right)}{\left(x^3-1\right)^2} \]

OpenStudy (anonymous):

24

OpenStudy (riley):

nah, definitely 27